Bob wants to have concrete driveway in front of his house. The driveway will measure 58 2/3 feet long and 18 ˝feet wide; and will be 3/4 foot thick. How many cubic yards of cement will be needed for the driveway? Round your answer to the nearest cubic yard. [1 cubic yd = 27 cubic feet]
Answer by christo(3)

V=L*W*H
V= 58.66*18.5*.75
= approx 814 cubic feet
which rounds out to 30 cubic yards, 814/27 = approx 30.